To find volume of water, which should be added to 11 of a 0,6 % solution of an acetic acid (p = 1) for reception of a solution with pH =3 (Kd = 1,8 10).
1
Expert's answer
2020-09-17T07:16:13-0400

The concentration of the solution is unknown.

The pH of the solution is "3.00"

The new volume is unknown.

The equilibrium equation is;

"CH_3COOH\leftrightharpoons H^++CH_3COO^-" At equilibrium, "C_o=1.8\times 10^-5"

"K_d=""[H^+][CH_3COO^-]\over[CH_3COOH]"

"=" "[1.8\times 10^{-5}]\times[1.8\times 10^{-5}]\over [-1.8\times 10^{-5}]" "=3.6\times 10^{-5}"

"[H^+]=10^{-3.00}"

The volume in litres of the solution is;

"V=" "1\over 3.6\times 10^{-5}" "=2.7\times 10^4" "L"

T

"0.6\over 100" "\times 2.7\times 10^4=1.6\times 10^2L"


The new volume "=" "1.6\times 10^2 L"